portfolio and investment analysis

A stockbroker calls you and suggests that you invest in the Lauren Computer Company. After analyzing the firm’s annual report and other material, you believe that the distribution of expected rates of return is as follows:


LAUREN COMPUTER CO.
Possible Rate of ReturnProbability
-0.10
0.25
0.05
0.20
0.15
0.25
0.25
0.10
0.40
0.10
0.55
0.10


Compute the expected return [E(Ri)] on Lauren Computer stock. Use a minus sign to enter a negative value, if any. Round your answer to one decimal place.

  %
Answer #1

Compute expected return \(\left[\mathrm{E}\left(\mathrm{R}_{\mathrm{i}}\right)\right]:\)

Formula for Expected Return is below:

Expected return \(=\sum_{i=1}^{n}\) (Probability of return)(Possible Return)

Calculate:

Expected return = (-0.10×0.25)+(0.05×0.20)+(0.15×0.25)+(0.25×0.10)+(0.40×0.10)+(0.55×0.10) = 0.1425 = 0.16 (or 16.00%)


Therefore, expected return \(\left[\mathrm{E}\left(\mathrm{R}_{\mathrm{i}}\right)\right]=\) 16.00%

answered by: Pushkin